Find the total resistance in each case:


a. A 16 Ω and an 18 Ω resistor connected in parallel

b. A 20 Ω, 10 Ω and 5 Ω resistor connected in parallel


1
Expert's answer
2021-12-28T09:37:41-0500

Solution;

For resistors in parallel;

"\frac{1}{R_T}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}+...+\frac{1}{R_n}"

(a)

"\frac{1}{R_T}=\frac{1}{16}+\frac{1}{18}=\frac{17}{144}"

Therefore;

"R_T=\frac{144}{17}=8.471\Omega"

(b)

"\frac{1}{R_T}=\frac{1}{20}+\frac{1}{10}+\frac{1}{5}=\frac{7}{20}"

Hence;

"R_T=2.857\Omega"